OpenStudy (anonymous):

Deshawn built a model train car. The length of the model train car is 6 in. The scale Deshawn used to build the model train car is 1:36. What is the actual length of the train car in FEET?

OpenStudy (anonymous):

\[3inches \times \frac{36inches}{inch}=198inches\]\[198inches \times \frac{1foot}{12inches}=16.5feet\] To check: \[0.5feet \times \frac{33inches}{inch}= 16.5feet\]